OpenStudy (anonymous):

there are 90 students in a speech contest yesterday 2/5 of them gave their speeches today 1/6 of the remaining students gave their speeches. how many students still havent givin their speeches

Parth (parthkohli):

40% + 16 2/3% + x(remaining) = 100% 16 2/3% + x = 60% x = 43 1/3% 43 1/3% of 90 = 39 39 is the answer.